if you have a family of 4 children what is the probability of having 1 boy?
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (sweetburger):
1/2 it does not matter how many children you have
OpenStudy (anonymous):
let me reword this.... you have a family of 4 children what is the probability that at least 1 is a boy?
OpenStudy (jonnyvonny):
Every 1 child, there was a chance to have either a boy or girl; so 1/2 per child, and that chance times itself 4 times, for you only have 1 boy out of 3.\[(1/2)*(1/2)*(1/2)*(1/2)= 1/8=.125=12.5\]
OpenStudy (jonnyvonny):
12.5%, wait, so you want ONLY 1 boy or at least 1?
OpenStudy (anonymous):
Is it one boy or at least one boy?..
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (anonymous):
at least one boy
OpenStudy (anonymous):
P(at least 1 boy)=1-P(all girls)=1-1/2*1/2*1/2*1/2
OpenStudy (anonymous):
At least one boy: 15/16
OpenStudy (anonymous):
how did you get 15/16
OpenStudy (anonymous):
Look at eashy's answer
Still Need Help?
Join the QuestionCove community and study together with friends!
Sign Up
OpenStudy (jonnyvonny):
Are we including hermaphadites?...
OpenStudy (anonymous):
Not sure if you knew stef but this is actually a famous mathematical paradox.. Look up the Boy or Girl paradox
OpenStudy (anonymous):
eashy did you get 15/16 also?
OpenStudy (anonymous):
Simply put the chance is 15/16 because the only chance that the children wouldn't include a boy is if they were all girls.
OpenStudy (anonymous):
P(at least one boy)+P(all girls)=1
Still Need Help?
Join the QuestionCove community and study together with friends!